### v1

# Hausman’s Consistency Test and an Internal Bias Diagnostic: A Short Note Regarding Linear Mixed Models

Abstract.
The Hausman specification test detects inconsistency in mixed model estimators of fixed effects by comparing the original model with an alternative specification in which random effects are treated as fixed. This note illustrates a bias diagnostic from the statistical literature as a complement to the Hausman test. The diagnostic can provide internal estimates of parameter-specific bias in the mixed model estimators without requiring the second, fixed-effects-only model to be fit. We apply the diagnostic to a panel data analysis as well as to a Value-Added Model (VAM) for teacher evaluation.
Key words and phrases:
linear mixed model, misspecification, permutation test, stochastic model matrix
1. Introduction
Consider a linear mixed model:
(1) ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon},$
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},{\mathbf{R}})$. For the resulting estimator of $\boldsymbol{\beta}$ to be consistent, it is necessary that the random effects $\boldsymbol{\eta}$ are uncorrelated with the regressors in ${\mathbf{X}}$ (Hausman & Taylor, 1981a). The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-$\boldsymbol{\eta}$ effects (RE) and fixed-$\boldsymbol{\eta}$ effects (FE) estimators, the latter requiring a re-fit of the model that treats $\boldsymbol{\eta}$ as an additional vector of fixed effects. Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them. Under the alternative, only the FE estimator is consistent. A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity. By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size. While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias. To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification.
We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when the random-effects design matrix ${\mathbf{Z}}$ may be correlated with the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based p-value, all without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability. For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test. While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between ${\mathbf{Z}}$ and $\boldsymbol{\eta}$. Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
In Section 2, we present both the Hausman test and the bias diagnostic. Section 3 applies these methods to a gasoline consumption dataset using the R packages plm and mixedbiastest, and Section 4 illustrates the bias diagnostic in a value-added modeling (VAM) context using the R package GPvam.
2. Hausman and Bias Diagnostics
2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting ${\mathbf{M}}^{-}$ represent a generalized inverse of ${\mathbf{M}}$, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})^{\prime}[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})]^{-}(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})$
is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})$ (Hausman & Taylor, 1981a). A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
2.2. Bias Diagnostic
The RE estimator is
$\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. Under the assumption that ${\mathbf{X}}$ is fixed and ${\mathbf{Z}}$ is stochastic, Karl & Zimmerman (2021) show
$\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$
where
$\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}$
and that this bias is zero if any of the following three conditions are true: (1) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (2) if $\boldsymbol{\eta}$ and ${\mathbf{Z}}$ are independent, or (3) if $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator). The third condition indicates that the FE estimator remains unbiased in the presence of model misspecification that biases the RE estimator. Thus the internal bias estimate often approximates the difference between the RE and FE estimators.
The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal bias estimate for the RE estimator. A vector permutation function $\boldsymbol{\pi}(\cdot)$ is applied repeatedly to calculate draws $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}})$ from a reference distribution to assess if the observed bias is unusually large under the null hypothesis that $\boldsymbol{\eta}$ is independent of all other model components. The bias p-value is calculated as the proportion of reference values that are larger in absolute value than that of the internal plug-in bias estimate. This test has been implemented in the R package mixedbiastest (Karl, 2024): the package accepts mixed model objects from the popular lme4 package (Bates et al., 2015) and produces bias diagnostics for each parameter of $\boldsymbol{\beta}$ as well as custom contrasts. mixedbiastest currently only accepts lme4 models with diagonal random effects covariance matrices $({\mathbf{G}})$, but this matches the structure of most panel data applications and could be generalized as described by Karl & Zimmerman (2021).
3. Gasoline Consumption Data Application
We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test.
1# Load required libraries
2library(plm)
3library(lme4)
4library(mixedbiastest)
5# Load the Gasoline dataset
6data("Gasoline", package = "plm")
7# Define the model formula
8form <- lgaspcar ~ lincomep + lrpmg + lcarpcap
9# Within (FE) model
10wi <- plm(form, data = Gasoline, model = "within")
11# Random effects model
12re <- plm(form, data = Gasoline, model = "random")
13# Hausman test
14htest <- phtest(wi, re)
15# Results
16print(htest)
Listing 1:  Hausman Test for the Gasoline Dataset
Hausman Test
chisq = 302.8, df = 3, p-value < 2.2e-16
The significant result rejects the null hypothesis that the RE estimator is consistent.
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test
1# mixed model
2random_model <- lmer(lgaspcar ~ lincomep + lrpmg + lcarpcap + (1 | country),
3 data = Gasoline)
4# run the bias test
5res<-mixedbiastest(random_model, n_permutations = 1e6, verbose = FALSE)
6# treat country indicators as fixed effects
7fixed_model <- lm(lgaspcar ~ lincomep + lrpmg + lcarpcap + country,
8 data = Gasoline)
9print(random_model)
10print(fixed_model)
11print(res)
Listing 2:  Bias Test for the Gasoline Dataset
Figure 1.  Graphical Summary of Bias Diagnostics for Gasoline Data. The histograms show the reference distributions, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}}),$ and the dashed red lines show the internal bias estimates, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$.
Table 1.  Gasoline Data: FE, RE Estimates, RE - FE difference, Internal Bias Estimates, and Bias p-value (rounded)
Parameter FE Est RE Est (RE - FE) Bias Estimate Bias p-value
(Intercept) 2.29 2.15 -0.14 -0.17 0.1048
lincomep 0.66 0.59 -0.07 -0.04 0.1596
lrpmg -0.32 -0.37 -0.05 -0.04 0.0008
lcarpcap -0.64 -0.62 0.02 0.01 0.1975
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test indicated inconsistency of the RE estimator, and bias diagnostic now suggests that lrpmg is significantly biased downward. Note the similarity of the internal bias estimates from the RE model with the RE-FE differences in Table 1.
4. Value-added Model (VAM) Results with GPvam
Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014). There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix (${\mathbf{R}}$). Classroom-heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$). A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix ${\mathbf{G}}$ with the number of unique elements equal to the number of grades in the study. The random effects model matrix ${\mathbf{Z}}$ tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple member” structure for ${\mathbf{Z}}$ (Browne et al., 2001). The nonrandom assignment of students to classrooms creates the potential for biasing (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021). Calculating the Hausman test statistic to test consistency for this example would require customized programming to fit a FE model that treats the teacher effects as fixed while still utilizing the multiple membership structure of ${\mathbf{Z}}$ and the block diagonal structure of the error covariance matrix, $\mathbf{R}$. By contrast, the bias test only requires the estimates and model matrices from the fitted mixed model.
The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic. Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the ${\mathbf{Z}}$, ${\mathbf{G}}$, and $\mathbf{R}$ matrices can be constructed. This is implicit in the function call, and the formula argument to GPvam only constructs the fixed effects structure ${\mathbf{X}}$ and $\boldsymbol{\beta}$. We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6. This is the same data set used by Karl et al. (2013a).
1library(GPvam)
2result.cp<-GPvam(data, persistence = "CP",
3 formula(~as.factor(Race_Ethnicity)+as.factor(year)+as.factor(Gender)+0),
4 hessian = TRUE,tol1=1e-9)
5# Perform the bias test on all fixed effects
6bias.test.custom(result.cp)
7# Perform the bias test on a contrast: White - Hispanic
8bias.test.custom(result.cp,n_perms=1e6, k_vectors = list(c(0,0,0,-1,1,0,0,0)))
Listing 3:  Fitting a CP model with GPvam and performing the bias diagnostic
Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Bias p-values
Fixed Effect Estimate Std. Error Bias Estimate Bias p-value
Black 24.2894 0.1612 0.0062 0.7430
Am Indian/Alaska 24.3683 0.3228 -0.0364 0.1570
Asian/Pac Island 25.7910 0.1863 0.0506 0.0105
Hispanic 24.3989 0.1058 -0.0691 0.0004
White 25.4836 0.0894 0.0597 0.0000
year=1 - - - -
year=2 0.9831 0.0600 0.0023 0.5440
year=3 1.9738 0.0876 0.0012 0.7530
Gender=F - - - -
Gender=M 0.0552 0.0819 0.0036 0.4990
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased. The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a p-value of 0 (none of the one million permutations exceeded the test statistic). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis.
5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their corresponding model matrix in situations where the fixed and random effects model matrices are not orthogonal (with respect to the inverse of the error covariance matrix). In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-$\boldsymbol{\eta}$ estimator of the fixed effects, provided the fixed-$\boldsymbol{\eta}$ estimator is consistent. Nevertheless, the bias diagnostic can rely solely on the fitted random-$\boldsymbol{\eta}$ model, generate summaries for each fixed effect or contrast, and use a permutation-based reference distribution. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.

### v2

@@
-Hausman’s Consistency Test and an Internal Bias Diagnostic: A Short Note Regarding Linear Mixed Models
+Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models
 Abstract.
-The Hausman specification test detects inconsistency in mixed model estimators of fixed effects by comparing the original model with an alternative specification in which random effects are treated as fixed.
-This note illustrates a bias diagnostic from the statistical literature as a complement to the Hausman test.
-The diagnostic can provide internal estimates of parameter-specific bias in the mixed model estimators without requiring the second, fixed-effects-only model to be fit.
-We apply the diagnostic to a panel data analysis as well as to a Value-Added Model (VAM) for teacher evaluation.
+The Hausman specification test detects inconsistency in random-effects estimators by comparing them with alternative fixed-effects estimators.
+This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications.
+The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model.
+We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
 Key words and phrases:
-linear mixed model, misspecification, permutation test, stochastic model matrix
+linear mixed model, Hausman test, correlated random effects, permutation test
 1. Introduction
@@
 with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},\mathbf{R})$.
-For the resulting estimator of $\boldsymbol{\beta}$ to be consistent, it is necessary that the random effects $\boldsymbol{\eta}$ are uncorrelated with the regressors in ${\mathbf{X}}$ (Hausman & Taylor, 1981a).
-The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-$\boldsymbol{\eta}$ effects (RE) and fixed-$\boldsymbol{\eta}$ effects (FE) estimators, the latter requiring a re-fit of the model that treats $\boldsymbol{\eta}$ as an additional vector of fixed effects.
+Here $\mathbf{y}$} is an $n\times 1$ response vector, $\mathbf{X}$} is an $n\times p$ fixed-effects design matrix with associated coefficient vector $\boldsymbol{\beta}$, and $\mathbf{Z}$} is an $n\times m$ random-effects design matrix with random-effects vector $\boldsymbol{\eta}$.
+The matrices $\mathbf{G}$} and $\mathbf{R}$ denote the $m\times m$ random-effects covariance matrix and the $n\times n$ error covariance matrix, respectively.
+In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors (the usual exogeneity condition; see, e.g., Hausman & Taylor, 1981a).
+The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-effects (RE) and fixed-effects (FE) estimators.
+In its original formulation, this involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests based on a single augmented RE fit; see Section 2.
 Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them.
@@
 To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification.
-We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when the random-effects design matrix ${\mathbf{Z}}$ may be correlated with the random effects $\boldsymbol{\eta}$.
+In what follows, we deliberately keep these roles separate: the Hausman test is used as a classical large-sample specification test for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to quantify the magnitude and direction of parameter-specific bias in the same RE fit.
+We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} may be stochastically dependent on the random effects $\boldsymbol{\eta}$.
 Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value, all without requiring a second, fixed-effects-only model fit.
@@
 Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
+Our aim is not to propose a new alternative to the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing linear mixed model bias diagnostic can supplement them by providing coefficient- or contrast-specific information obtained from a single fitted mixed model.
 In Section 2, we present both the Hausman test and the bias diagnostic.
@@
 Under the null hypothesis, both estimators are consistent.
-Letting ${\mathbf{M}}^{-}$ represent a generalized inverse of ${\mathbf{M}}$, the test statistic
-$H=(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})^{\prime}[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})]^{-}(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})$
-is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})$ (Hausman & Taylor, 1981a).
+Letting $\mathbf{M}$$^{-}$ represent a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
+$H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
+is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a).
 A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
+In the canonical panel-data setting, this statistic is derived under the assumption of homoskedastic, serially uncorrelated errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$.
+In the linear mixed model formulation we write the model with a general error covariance matrix $\mathbf{R}$; the gasoline example below maintains the standard homoskedastic assumption, whereas the value-added model in Section 4 uses a block-diagonal $\mathbf{R}$.
 2.2. Bias Diagnostic
-The RE estimator is
-$\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
+The random-effects (RE) estimator of $\boldsymbol{\beta}$ is
+(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
 with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$.
-Under the assumption that ${\mathbf{X}}$ is fixed and ${\mathbf{Z}}$ is stochastic, Karl & Zimmerman (2021) show
-$\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$
-where
-$\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}$
-and that this bias is zero if any of the following three conditions are true: (1) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (2) if $\boldsymbol{\eta}$ and ${\mathbf{Z}}$ are independent, or (3) if $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator).
-The third condition indicates that the FE estimator remains unbiased in the presence of model misspecification that biases the RE estimator.
-Thus the internal bias estimate often approximates the difference between the RE and FE estimators.
-The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal bias estimate for the RE estimator.
-A vector permutation function $\boldsymbol{\pi}(\cdot)$ is applied repeatedly to calculate draws $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}})$ from a reference distribution to assess if the observed bias is unusually large under the null hypothesis that $\boldsymbol{\eta}$ is independent of all other model components.
-The bias p-value is calculated as the proportion of reference values that are larger in absolute value than that of the internal plug-in bias estimate.
-This test has been implemented in the R package mixedbiastest (Karl, 2024): the package accepts mixed model objects from the popular lme4 package (Bates et al., 2015) and produces bias diagnostics for each parameter of $\boldsymbol{\beta}$ as well as custom contrasts.
-mixedbiastest currently only accepts lme4 models with diagonal random effects covariance matrices $({\mathbf{G}})$, but this matches the structure of most panel data applications and could be generalized as described by Karl & Zimmerman (2021).
+Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic.
+Let ${\mathbf{V}}={\mathbf{Z}}{\mathbf{G}}{\mathbf{Z}}^{\prime}+\mathbf{R}$ denote the true marginal covariance matrix of $\mathbf{y}$}, and define
+$\boldsymbol{\nu_{k}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{Z}}.$
+Karl & Zimmerman (2021) show that for any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$),
+(3) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\boldsymbol{\nu_{k}}^{\prime}\boldsymbol{\eta}].$
+Karl & Zimmerman (2021) propose a plug-in approximation based on the fitted mixed model. Writing
+(4) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}},$
+and denoting the empirical best linear unbiased predictors (EBLUPs) of the random effects by $\hat{\boldsymbol{\eta}}$, the plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal estimate of the bias of $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$.
+The exact bias is zero if any of the following three conditions hold: (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, or (iii) $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator).
+The third condition implies that the FE estimator remains unbiased in settings where the RE estimator is biased, so that the internal bias estimate often approximates the difference between the RE and FE estimators (Karl & Zimmerman, 2021).
+[…]
+2.3. Relation to correlated random effects and Mundlak–Wooldridge tests
+In the panel-data literature, a common way to test the random-effects specification is through correlated random effects (CRE) or “Mundlak” type augmentations of the RE model.
+Suppose we have the panel-data setup
+$y_{it}=x_{it}^{\prime}\beta+c_{i}+\varepsilon_{it},$
+with individual effect $c_{i}$ and idiosyncratic error $\varepsilon_{it}$.
+Under the usual random-effects assumptions, including strict exogeneity and $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ (often summarized as $\operatorname{Cov}(x_{it},c_{i})=0$), the random-effects estimator is consistent.
+Mundlak (1978) models potential correlation between $x_{it}$ and $c_{i}$ by including the group means $\bar{x}_{i}$ in the regression.
+Wooldridge (2010) shows that a simple Hausman-type alternative is to estimate the augmented random-effects model
+$y_{it}=x_{it}^{\prime}\beta+\bar{x}_{i}^{\prime}\gamma+c_{i}+\varepsilon_{it},$
+and apply a (robust) Wald test of $H_{0}:\gamma=0$.
+This CRE test uses only an augmented RE fit but is asymptotically equivalent to the traditional FE–RE Hausman test when the idiosyncratic errors are homoskedastic and serially uncorrelated (Hausman & Taylor, 1981a).
+Robust implementations based on covariance estimators that are consistent under general forms of heteroskedasticity and serial correlation are widely used in applied work; see, e.g., Wooldridge (2010).
+In terms of the linear mixed model (1), stacking the panel outcomes $\{y_{it}\}$ into $\mathbf{y}$} and the regressors $\{x_{it}\}$ into the rows of $\mathbf{X}$} yields ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon}$, where $\boldsymbol{\eta}$ collects the individual effects $c_{i}$ and $\mathbf{Z}$} is the $n\times N$ incidence matrix (with $N$ the number of individuals) that has a one in the row corresponding to observation $(i,t)$ and column $i$ when that observation belongs to individual $i$.
+In this formulation the assignment structure encoded by $\mathbf{Z}$} may itself be stochastic and potentially correlated with the random effects $\boldsymbol{\eta}$, which is precisely the type of dependence considered by Karl & Zimmerman (2021).
+In the conventional RE panel-data model, the exogeneity condition $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ restricts the dependence between the random effects $c_{i}$ and the regressors $x_{it}$.
+By contrast, the zero-bias conditions listed in the previous subsection include a requirement that rules out dependence between the random effects and the random-effects design matrix $\mathbf{Z}$}.
+In simple random-intercept panels $\mathbf{Z}$} is just an incidence matrix and is often implicitly assumed to be fixed without further consideration.
+Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that arises when the random-effects design matrix $\mathbf{Z}$} and the random-effects vector $\boldsymbol{\eta}$ are dependent, with $\mathbf{X}$} treated as fixed.
+The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ and its permutation-based $p$-value can therefore be viewed as a parameter-specific analogue of a Hausman-type specification check, obtained from a single fitted mixed model.
+In applications where both $\mathbf{X}$} and $\mathbf{Z}$} may reasonably be regarded as stochastic, the diagnostic still targets the component of bias that operates through the joint behavior of $\mathbf{Z}$} and $\boldsymbol{\eta}$; it does not attempt to separate any additional bias that might arise from stochastic variation in $\mathbf{X}$} alone.
+In this sense, the Karl & Zimmerman (2021) diagnostic focuses on a narrower class of misspecification than the classical Hausman or Mundlak–Wooldridge tests, and is best viewed as a complementary, parameter-specific tool rather than a replacement for those global specification checks.
+An advantage of this approach is that it can be implemented directly from a single fitted mixed model, without requiring the analyst to construct a full FE or CRE analogue in settings that depart from the standard individual-by-time panel, such as models with multiple-membership random effects, nontrivial error-covariance structures $\mathbf{R}$, or a very large number of random-effect levels.
+In such applications, FE or CRE-style specifications are often still possible in principle, but may require additional data construction and modeling work.
+The value-added model in Section 4 provides one such example.
+When conventional FE or CRE estimators are readily available, we recommend using the classical Hausman or Mundlak–Wooldridge tests as global checks of the RE specification, and then using the bias diagnostic to identify which coefficients or contrasts exhibit the largest finite-sample bias and in which direction.
 3. Gasoline Consumption Data Application
@@
 Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test.
+All results in this section can be reproduced using Listings 1 and 2 together with the Gasoline data supplied by the plm package.
 1# Load required libraries
@@
 Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2).
-For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test
+For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test.
 1# mixed model
@@
 Listing 2:  Bias Test for the Gasoline Dataset
-Figure 1.
-Graphical Summary of Bias Diagnostics for Gasoline Data.
-The histograms show the reference distributions, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}}),$ and the dashed red lines show the internal bias estimates, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$.
+[…]
 Table 1.  Gasoline Data: FE, RE Estimates, RE - FE difference, Internal Bias Estimates, and Bias p-value (rounded)
@@
 The results from Listing 2 are combined into Table 1.
-The Hausman test indicated inconsistency of the RE estimator, and bias diagnostic now suggests that lrpmg is significantly biased downward.
-Note the similarity of the internal bias estimates from the RE model with the RE-FE differences in Table 1.
+The Hausman test indicates global inconsistency of the RE estimators, and the bias diagnostic suggests that the RE estimate of lrpmg is significantly biased downward.
+Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
 4. Value-added Model (VAM) Results with GPvam
@@
 In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix ($\mathbf{R}$).
-Classroom-heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$).
+Classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$).
 A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with the number of unique elements equal to the number of grades in the study.
-The random effects model matrix ${\mathbf{Z}}$ tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple member” structure for ${\mathbf{Z}}$ (Browne et al., 2001).
-The nonrandom assignment of students to classrooms creates the potential for biasing (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021).
-Calculating the Hausman test statistic to test consistency for this example would require customized programming to fit a FE model that treats the teacher effects as fixed while still utilizing the multiple membership structure of ${\mathbf{Z}}$ and the block diagonal structure of the error covariance matrix, $\mathbf{R}$.
-By contrast, the bias test only requires the estimates and model matrices from the fitted mixed model.
+The random-effects model matrix $\mathbf{Z}$} tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple-membership” structure for $\mathbf{Z}$} (Browne et al., 2001).
+The nonrandom assignment of students to classrooms creates the potential for bias (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021), through dependence between the random-effects design matrix $\mathbf{Z}$} and the teacher effects $\boldsymbol{\eta}$.
+These features make the CP VAM a natural setting in which to use a parameter-specific bias diagnostic.
+The random-effects structure involves thousands of teacher effects, the student-level error covariance matrix $\mathbf{R}$ is block-diagonal rather than homoskedastic, and the multiple-membership design encoded by $\mathbf{Z}$} is more complex than a simple random-intercept panel.
+The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
 The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic.
 Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices can be constructed.
-This is implicit in the function call, and the formula argument to GPvam only constructs the fixed effects structure ${\mathbf{X}}$ and $\boldsymbol{\beta}$.
+This is implicit in the function call, and the formula argument to GPvam only constructs the fixed-effects design matrix $\mathbf{X}$} and parameter vector $\boldsymbol{\beta}$.
 We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6.
@@
 (2013a).
+In this value-added model application, we use the GPvam implementation of the bias diagnostic, which accommodates the block-diagonal error covariance matrix and multiple-membership random effects structure that fall outside the scope of the current mixedbiastest implementation.
 1library(GPvam)
@@
 Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis.
+In this example we do not propose a change in estimator based solely on the diagnostic.
+Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.
 5. Conclusion
 We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects.
-Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their corresponding model matrix in situations where the fixed and random effects model matrices are not orthogonal (with respect to the inverse of the error covariance matrix).
-In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-$\boldsymbol{\eta}$ estimator of the fixed effects, provided the fixed-$\boldsymbol{\eta}$ estimator is consistent.
-Nevertheless, the bias diagnostic can rely solely on the fitted random-$\boldsymbol{\eta}$ model, generate summaries for each fixed effect or contrast, and use a permutation-based reference distribution.
+Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their design matrix $\mathbf{Z}$} in situations where the fixed- and random-effects design matrices are not orthogonal (with respect to the inverse of the error covariance matrix).
+In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent.
+Nevertheless, the bias diagnostic can be computed from the fitted random-effects model without changing the original RE specification and generates summaries for each fixed effect or contrast.
 These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.
+In applications where both RE and FE (or CRE) estimators are readily available, we view the Hausman or Mundlak–Wooldridge tests as primary specification checks: a rejection of the null typically leads the analyst to prefer an estimator that is robust to correlation between regressors and random effects.
+The bias diagnostic is then informative about which coefficients or contrasts are most affected and by how much.
+When the conclusions from the two approaches differ, we recommend using the diagnostic in a sensitivity-analysis role rather than as a replacement for the classical specification tests.
+A convenient workflow is therefore to (i) fit the RE model, (ii) apply a FE- or CRE-based Hausman/Mundlak–Wooldridge test, and (iii) when the test is rejected or borderline, use the mixed-model bias diagnostic to identify and report the coefficients or contrasts that are most affected.

### v3

@@
 Abstract.
-The Hausman specification test detects inconsistency in random-effects estimators by comparing them with alternative fixed-effects estimators.
+The Hausman specification test detects inconsistency of the random-effects estimator by comparing it with an alternative fixed-effects estimator.
 This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications.
-The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model.
+The diagnostic delivers parameter-specific internal estimates of finite-sample bias of the random-effects estimator, together with permutation-based $p$-values, from a single fitted random-effects model.
 We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
@@
 2.2. Bias Diagnostic
-The random-effects (RE) estimator of $\boldsymbol{\beta}$ is
-(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
-with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$.
+Let $\hat{{\mathbf{G}}}$ and $\hat{\mathbf{R}}$ denote estimates of $\mathbf{G}$} and $\mathbf{R}$, and define $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$.
+The empirical random-effects (RE) estimator of $\boldsymbol{\beta}$ is
+(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}}.$
 Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic.
-Let ${\mathbf{V}}={\mathbf{Z}}{\mathbf{G}}{\mathbf{Z}}^{\prime}+\mathbf{R}$ denote the true marginal covariance matrix of $\mathbf{y}$}, and define
-$\boldsymbol{\nu_{k}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{Z}}.$
-Karl & Zimmerman (2021) show that for any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$),
-(3) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\boldsymbol{\nu_{k}}^{\prime}\boldsymbol{\eta}].$
-Karl & Zimmerman (2021) propose a plug-in approximation based on the fitted mixed model. Writing
-(4) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}},$
-and denoting the empirical best linear unbiased predictors (EBLUPs) of the random effects by $\hat{\boldsymbol{\eta}}$, the plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal estimate of the bias of $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$.
-The exact bias is zero if any of the following three conditions hold: (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, or (iii) $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator).
-The third condition implies that the FE estimator remains unbiased in settings where the RE estimator is biased, so that the internal bias estimate often approximates the difference between the RE and FE estimators (Karl & Zimmerman, 2021).
+For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define
+(3) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
+Karl & Zimmerman (2021) show that, under regularity conditions,
+(4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
+The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ or (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}.
+A third way to avoid bias in practice is to treat $\boldsymbol{\eta}$ as fixed (producing the FE estimator), in which case the corresponding fixed-effects estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased.
 […]
@@
 For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test.
+Although the random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, they are very similar in this simple random-intercept setting.
+We therefore use plm for the Hausman test and lmer for the bias diagnostic, since mixedbiastest is implemented on top of lme4.
 1# mixed model
@@
 The results from Listing 2 are combined into Table 1.
-The Hausman test indicates global inconsistency of the RE estimators, and the bias diagnostic suggests that the RE estimate of lrpmg is significantly biased downward.
+The Hausman test rejects the null hypothesis that the RE estimator is consistent, and the bias diagnostic suggests that the RE estimator for lrpmg is downward biased.
 Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
@@
 The CP estimates and bias diagnostics for the fixed effects appear in Table 2.
-The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased.
-The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a $p$-value of 0 (none of the one million permutations exceeded the test statistic).
+The diagnostic indicates that the estimator for the Hispanic effect is downward biased, while the estimators for the White and Asian/Pac Island effects are upward biased.
+The diagnostic for the contrast White - Hispanic reveals even stronger bias: the estimated contrast is $0.1287$, and none of the one million permutations exceeded the test statistic ($p$-value $\approx 0$).
 Although the bias is statistically significant, its magnitude may not be practically significant for this particular application.

### v4

@@
 In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test.
-While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between $\mathbf{Z}$} and $\boldsymbol{\eta}$.
+While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic summarizes finite-sample bias through the quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, which is sensitive to dependence between the random effects and the model-dependent weighting vector $\hat{\boldsymbol{\nu_{k}}}$ (and, in the stochastic-$\mathbf{Z}$} setting emphasized by Karl & Zimmerman, 2021, this dependence is induced by nonrandom $\mathbf{Z}$} that is related to $\boldsymbol{\eta}$).
 Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
@@
 $H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
-is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a).
+is $\chi^{2}(r)$ under $H_{0}$, where $r$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a).
 A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
@@
 Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic.
-For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define
-(3) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
+For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define the $1\times m$ row vector
+(3) $\hat{\boldsymbol{\nu_{k}}}^{\prime}\equiv\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
 Karl & Zimmerman (2021) show that, under regularity conditions,
 (4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
-The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ or (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}.
-A third way to avoid bias in practice is to treat $\boldsymbol{\eta}$ as fixed (producing the FE estimator), in which case the corresponding fixed-effects estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased.
+The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ (orthogonality with respect to $\mathbf{R}^{-1}$, with probability one), or (ii) $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]=0$ (which can be interpreted as the random effects being exogenous with respect to the model-dependent weighting embodied by $\hat{\boldsymbol{\nu_{k}}}$).
+In the stochastic-$\mathbf{Z}$} framework emphasized by Karl & Zimmerman (2021), with $\mathbf{X}$} treated as fixed, a sufficient condition for (ii) is that $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}.
+A third way to avoid this source of bias is to treat $\boldsymbol{\eta}$ as fixed (i.e., fit an FE model), in which case the corresponding FE estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased under the same $\boldsymbol{\epsilon}\perp{\mathbf{Z}}$ condition (Karl & Zimmerman, 2021).
 […]
@@
 In simple random-intercept panels $\mathbf{Z}$} is just an incidence matrix and is often implicitly assumed to be fixed without further consideration.
-Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that arises when the random-effects design matrix $\mathbf{Z}$} and the random-effects vector $\boldsymbol{\eta}$ are dependent, with $\mathbf{X}$} treated as fixed.
+Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that operates through the dependence captured by $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$ (in Karl & Zimmerman, 2021, this is induced by dependence between stochastic $\mathbf{Z}$} and $\boldsymbol{\eta}$, with $\mathbf{X}$} treated as fixed).
 The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ and its permutation-based $p$-value can therefore be viewed as a parameter-specific analogue of a Hausman-type specification check, obtained from a single fitted mixed model.
@@
 lcarpcap -0.64 -0.62 0.02 0.01 0.1975
+Remark.
+The “FE Est” intercept in Table 1 comes from the least-squares fit with country dummies and therefore depends on the dummy-variable normalization (choice of baseline category).
+It is not directly comparable to the intercept parameter in the random-intercept model, so we focus on the slope coefficients when interpreting FE–RE differences.
 The graphical summaries of the bias estimates and reference distributions appear in Figure 1.
@@
 We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects.
-Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their design matrix $\mathbf{Z}$} in situations where the fixed- and random-effects design matrices are not orthogonal (with respect to the inverse of the error covariance matrix).
+Besides considering bias rather than consistency, this diagnostic is narrower than the Hausman test in the sense that it targets the component of bias captured by $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$, i.e., bias operating through systematic alignment of the random effects with the weighting vector $\hat{\boldsymbol{\nu_{k}}}$ when the fixed- and random-effects design matrices are not orthogonal (with respect to $\mathbf{R}^{-1}$).
+In the stochastic-$\mathbf{Z}$} setting emphasized by Karl & Zimmerman (2021), such alignment is induced by dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$.
 In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent.

### v5

@@
 Abstract.
-The Hausman specification test detects inconsistency of the random-effects estimator by comparing it with an alternative fixed-effects estimator.
-This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications.
-The diagnostic delivers parameter-specific internal estimates of finite-sample bias of the random-effects estimator, together with permutation-based $p$-values, from a single fitted random-effects model.
-We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
+The Hausman specification test assesses the random-effects specification by comparing the random-effects estimator with a fixed-effects alternative.
+This note shows how a recently proposed bias diagnostic for linear mixed models can complement that test in random-effects panel-data applications.
+The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model.
+We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation using publicly available R packages, and we discuss how the resulting coefficient-specific bias summaries can be incorporated into routine practice.
 Key words and phrases:
@@
 The matrices $\mathbf{G}$} and $\mathbf{R}$ denote the $m\times m$ random-effects covariance matrix and the $n\times n$ error covariance matrix, respectively.
-In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors (the usual exogeneity condition; see, e.g., Hausman & Taylor, 1981a).
-The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-effects (RE) and fixed-effects (FE) estimators.
-In its original formulation, this involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests based on a single augmented RE fit; see Section 2.
-Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them.
-Under the alternative, only the FE estimator is consistent.
-A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
-The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity.
-By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size.
-While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias.
-To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification.
-In what follows, we deliberately keep these roles separate: the Hausman test is used as a classical large-sample specification test for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to quantify the magnitude and direction of parameter-specific bias in the same RE fit.
-We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} may be stochastically dependent on the random effects $\boldsymbol{\eta}$.
-Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value, all without requiring a second, fixed-effects-only model fit.
-This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability.
-For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
-In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test.
-While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic summarizes finite-sample bias through the quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, which is sensitive to dependence between the random effects and the model-dependent weighting vector $\hat{\boldsymbol{\nu_{k}}}$ (and, in the stochastic-$\mathbf{Z}$} setting emphasized by Karl & Zimmerman, 2021, this dependence is induced by nonrandom $\mathbf{Z}$} that is related to $\boldsymbol{\eta}$).
-Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
-Our aim is not to propose a new alternative to the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing linear mixed model bias diagnostic can supplement them by providing coefficient- or contrast-specific information obtained from a single fitted mixed model.
-In Section 2, we present both the Hausman test and the bias diagnostic.
-Section 3 applies these methods to a gasoline consumption dataset using the R packages plm and mixedbiastest, and Section 4 illustrates the bias diagnostic in a value-added modeling (VAM) context using the R package GPvam.
+In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors; see, for example, Hausman & Taylor, 1981a.
+The Hausman test (Hausman, 1978) is the classical tool for assessing this specification by comparing the random-effects (RE) and fixed-effects (FE) estimators.
+In its original formulation, this comparison involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests from a single augmented RE fit; see Section 2.
+Under the null hypothesis, both RE and FE are consistent, so their estimates should differ only by sampling variation.
+Under the alternative, only the FE estimator remains consistent.
+A large discrepancy is therefore evidence against the RE specification (Hausman, 1978; Hausman & Taylor, 1981b).
+The Hausman test concerns consistency: it asks whether the RE estimator converges to the true parameter values as the sample size grows.
+Bias, by contrast, is the systematic difference between an estimator’s expectation and the target parameter at a given sample size.
+These are distinct concepts – an estimator may be asymptotically consistent yet still exhibit finite-sample bias – but similar forms of misspecification, especially failures of exogeneity, can affect both.
+It is therefore natural to use a bias diagnostic alongside the Hausman test when assessing possible misspecification.
+In what follows, we keep these roles separate: the Hausman test serves as a classical large-sample specification check for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to summarize the magnitude and direction of parameter-specific bias within the same RE fit.
+We draw on the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} is stochastic and may be dependent on the random effects $\boldsymbol{\eta}$.
+Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value without requiring a second, fixed-effects-only model fit.
+This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical because of insufficient degrees of freedom or the absence of off-the-shelf software.
+For the theoretical development of the bias formula and the permutation scheme, see Karl & Zimmerman (2021).
+In this note, we show how the bias diagnostic can be used alongside the Hausman specification test.
+The Hausman test is global and asymptotic, whereas the internal bias diagnostic is parameter-specific and finite-sample.
+The quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ is large when the fitted random effects align strongly with the model-dependent weighting vector $\hat{\boldsymbol{\nu_{k}}}$; in the stochastic-$\mathbf{Z}$} setting of Karl & Zimmerman, 2021, that alignment is induced by dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$.
+Our aim is not to propose a replacement for the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing mixed-model bias diagnostic can supplement them by providing coefficient- or contrast-specific information from a single fitted mixed model.
+Section 2 reviews the Hausman test and the bias diagnostic.
+Section 3 applies these ideas to a gasoline-consumption panel using the plm and mixedbiastest packages, and Section 4 illustrates the diagnostic in a value-added modeling (VAM) application using GPvam.
 2. Hausman and Bias Diagnostics
@@
 Under the null hypothesis, both estimators are consistent.
-Letting $\mathbf{M}$$^{-}$ represent a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
+Letting $\mathbf{M}$$^{-}$ denote a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
 $H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
-is $\chi^{2}(r)$ under $H_{0}$, where $r$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a).
-A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
-In the canonical panel-data setting, this statistic is derived under the assumption of homoskedastic, serially uncorrelated errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$.
+has an asymptotic $\chi^{2}(r)$ null distribution, where $r$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a).
+A significant value of $H$ provides evidence against the RE orthogonality assumptions and, under the usual conditions, against consistency of the RE estimator.
+In the canonical panel-data setting, this statistic is derived under homoskedastic, serially uncorrelated idiosyncratic errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$.
 In the linear mixed model formulation we write the model with a general error covariance matrix $\mathbf{R}$; the gasoline example below maintains the standard homoskedastic assumption, whereas the value-added model in Section 4 uses a block-diagonal $\mathbf{R}$.
@@
 (2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}}.$
-Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic.
-For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define the $1\times m$ row vector
+To connect directly with Karl & Zimmerman (2021), we first present the diagnostic under their setup, which conditions on $\mathbf{X}$} and treats $\mathbf{Z}$} as stochastic.
+For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define the $1\times m$ row vector
 (3) $\hat{\boldsymbol{\nu_{k}}}^{\prime}\equiv\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
-Karl & Zimmerman (2021) show that, under regularity conditions,
+Under the conditions of Theorem 2 of Karl & Zimmerman (2021),
 (4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
-The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ (orthogonality with respect to $\mathbf{R}^{-1}$, with probability one), or (ii) $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]=0$ (which can be interpreted as the random effects being exogenous with respect to the model-dependent weighting embodied by $\hat{\boldsymbol{\nu_{k}}}$).
-In the stochastic-$\mathbf{Z}$} framework emphasized by Karl & Zimmerman (2021), with $\mathbf{X}$} treated as fixed, a sufficient condition for (ii) is that $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}.
-A third way to avoid this source of bias is to treat $\boldsymbol{\eta}$ as fixed (i.e., fit an FE model), in which case the corresponding FE estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased under the same $\boldsymbol{\epsilon}\perp{\mathbf{Z}}$ condition (Karl & Zimmerman, 2021).
+Thus the finite-sample bias of the RE estimator for $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is governed by the alignment between the random effects $\boldsymbol{\eta}$ and the coefficient-specific weighting vector $\hat{\boldsymbol{\nu_{k}}}$.
+Plugging the empirical best linear unbiased predictor $\hat{\boldsymbol{\eta}}$ into the right-hand side of (4) yields the internal bias estimate $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$.
+This bias is zero, for example, if (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ with probability one, which implies $\hat{\boldsymbol{\nu_{k}}}=\boldsymbol{0}$, or (ii) in the stochastic-$\mathbf{Z}$} framework of Karl & Zimmerman (2021), $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, which implies $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]=0$.
+A third way to remove this source of bias is to treat $\boldsymbol{\eta}$ as fixed (that is, fit an FE model), in which case the corresponding FE estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased under the same $\boldsymbol{\epsilon}\perp{\mathbf{Z}}$ condition, provided that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains estimable (Karl & Zimmerman, 2021).
 […]
@@
 In terms of the linear mixed model (1), stacking the panel outcomes $\{y_{it}\}$ into $\mathbf{y}$} and the regressors $\{x_{it}\}$ into the rows of $\mathbf{X}$} yields ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon}$, where $\boldsymbol{\eta}$ collects the individual effects $c_{i}$ and $\mathbf{Z}$} is the $n\times N$ incidence matrix (with $N$ the number of individuals) that has a one in the row corresponding to observation $(i,t)$ and column $i$ when that observation belongs to individual $i$.
-In this formulation the assignment structure encoded by $\mathbf{Z}$} may itself be stochastic and potentially correlated with the random effects $\boldsymbol{\eta}$, which is precisely the type of dependence considered by Karl & Zimmerman (2021).
-In the conventional RE panel-data model, the exogeneity condition $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ restricts the dependence between the random effects $c_{i}$ and the regressors $x_{it}$.
-By contrast, the zero-bias conditions listed in the previous subsection include a requirement that rules out dependence between the random effects and the random-effects design matrix $\mathbf{Z}$}.
-In simple random-intercept panels $\mathbf{Z}$} is just an incidence matrix and is often implicitly assumed to be fixed without further consideration.
-Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that operates through the dependence captured by $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$ (in Karl & Zimmerman, 2021, this is induced by dependence between stochastic $\mathbf{Z}$} and $\boldsymbol{\eta}$, with $\mathbf{X}$} treated as fixed).
-The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ and its permutation-based $p$-value can therefore be viewed as a parameter-specific analogue of a Hausman-type specification check, obtained from a single fitted mixed model.
-In applications where both $\mathbf{X}$} and $\mathbf{Z}$} may reasonably be regarded as stochastic, the diagnostic still targets the component of bias that operates through the joint behavior of $\mathbf{Z}$} and $\boldsymbol{\eta}$; it does not attempt to separate any additional bias that might arise from stochastic variation in $\mathbf{X}$} alone.
-In this sense, the Karl & Zimmerman (2021) diagnostic focuses on a narrower class of misspecification than the classical Hausman or Mundlak–Wooldridge tests, and is best viewed as a complementary, parameter-specific tool rather than a replacement for those global specification checks.
+In a conventional random-intercept panel, this incidence matrix is usually treated as fixed once the sample has been defined.
+The classical RE exogeneity condition then concerns dependence between the random effects $c_{i}$ and the regressors $x_{it}$, whereas the formal stochastic-$\mathbf{Z}$} framework of Karl & Zimmerman (2021) concerns dependence between $\boldsymbol{\eta}$ and the assignment structure encoded by $\mathbf{Z}$}.
+The two setups are therefore related but not identical.
+For standard panels, we view $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ primarily as a descriptive, coefficient-specific summary from the fitted RE model rather than as a literal test of a stochastic-$\mathbf{Z}$} assignment mechanism.
+Because $\hat{\boldsymbol{\nu_{k}}}$ depends on the fitted RE weighting for the coefficient or contrast of interest, large values of $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ often accompany sizeable FE–RE differences for the same coefficient.
+In applications where the assignment structure encoded by $\mathbf{Z}$} is itself plausibly stochastic – for example, multiple-membership or assignment problems – the interpretation developed by Karl & Zimmerman (2021) applies more directly.
+In either case, the diagnostic is best viewed as complementary to, rather than a substitute for, the Hausman or Mundlak–Wooldridge tests: it does not test the global null $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$, but instead supplies a parameter-specific finite-sample summary from a single fitted mixed model.
 An advantage of this approach is that it can be implemented directly from a single fitted mixed model, without requiring the analyst to construct a full FE or CRE analogue in settings that depart from the standard individual-by-time panel, such as models with multiple-membership random effects, nontrivial error-covariance structures $\mathbf{R}$, or a very large number of random-effect levels.
@@
 3. Gasoline Consumption Data Application
-We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008).
-Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test.
-All results in this section can be reproduced using Listings 1 and 2 together with the Gasoline data supplied by the plm package.
+We illustrate both diagnostics using the Gasoline dataset from the plm package, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008).
+Listing 1 shows the code for the Hausman test, and Listing 2 shows the code for the bias diagnostic.
+All results in this section can be reproduced using those two listings together with the Gasoline data supplied by plm.
 1# Load required libraries
@@
 chisq = 302.8, df = 3, p-value < 2.2e-16
-The significant result rejects the null hypothesis that the RE estimator is consistent.
+This overwhelmingly small $p$-value provides strong evidence against the RE specification for the slope coefficients.
 Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2).
-For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test.
-Although the random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, they are very similar in this simple random-intercept setting.
+For comparison of coefficient estimates, we also fit the corresponding least-squares model with country dummies, although that second fit is not required by the diagnostic itself.
+The random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, but they are very close in this simple random-intercept setting.
 We therefore use plm for the Hausman test and lmer for the bias diagnostic, since mixedbiastest is implemented on top of lme4.
@@
 3 data = Gasoline)
-4# run the bias test
+4# run the bias diagnostic
 5res<-mixedbiastest(random_model, n_permutations = 1e6, verbose = FALSE)
@@
 11print(res)
-Listing 2:  Bias Test for the Gasoline Dataset
+Listing 2:  Bias Diagnostic for the Gasoline Dataset
 […]
-Table 1.  Gasoline Data: FE, RE Estimates, RE - FE difference, Internal Bias Estimates, and Bias p-value (rounded)
-Parameter FE Est RE Est (RE - FE) Bias Estimate Bias p-value
+Table 1.  Gasoline Data: FE and RE Estimates, RE - FE Difference, Internal Bias Estimates, and Permutation $p$-values (rounded)
+Parameter FE Est RE Est (RE - FE) Bias Estimate Perm. $p$-value
 (Intercept) 2.29 2.15 -0.14 -0.17 0.1048
@@
 It is not directly comparable to the intercept parameter in the random-intercept model, so we focus on the slope coefficients when interpreting FE–RE differences.
-The graphical summaries of the bias estimates and reference distributions appear in Figure 1.
-The results from Listing 2 are combined into Table 1.
-The Hausman test rejects the null hypothesis that the RE estimator is consistent, and the bias diagnostic suggests that the RE estimator for lrpmg is downward biased.
-Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
+The graphical summaries of the internal bias estimates and reference distributions appear in Figure 1, and the numerical results from Listing 2 are summarized in Table 1.
+Because the country-incidence matrix is fixed once the sample is defined, we interpret these bias summaries descriptively rather than as evidence for a literal stochastic-$\mathbf{Z}$} assignment mechanism.
+On that descriptive reading, the diagnostic points most clearly to lrpmg: its internal bias estimate is negative and its permutation $p$-value is very small.
+Note also the similarity between the internal bias estimates from the RE model and the corresponding FE–RE differences in Table 1.
 4. Value-added Model (VAM) Results with GPvam
-Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014).
-There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b).
-In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix ($\mathbf{R}$).
-Classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$).
-A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with the number of unique elements equal to the number of grades in the study.
-The random-effects model matrix $\mathbf{Z}$} tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple-membership” structure for $\mathbf{Z}$} (Browne et al., 2001).
-The nonrandom assignment of students to classrooms creates the potential for bias (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021), through dependence between the random-effects design matrix $\mathbf{Z}$} and the teacher effects $\boldsymbol{\eta}$.
-These features make the CP VAM a natural setting in which to use a parameter-specific bias diagnostic.
+Linear mixed models are widely used in value-added assessment of teachers based on their students’ standardized test scores (Morganstein & Wasserstein, 2014).
+One important instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b).
+In the CP model, intra-student correlation is represented through off-diagonal entries in the error covariance matrix $\mathbf{R}$.
+Between-classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled through random effects $\boldsymbol{\eta}$.
+A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with one distinct diagonal element for each grade in the study.
+The random-effects model matrix $\mathbf{Z}$} records students’ classroom assignments and also attributes post-baseline scores to prior teachers, producing a multiple-membership structure for $\mathbf{Z}$} (Browne et al., 2001).
+Nonrandom assignment of students to classrooms can therefore create the sort of dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$ highlighted by Karl & Zimmerman (2021), and hence the potential for bias (Lockwood & McCaffrey, 2007).
+These features make the CP VAM a natural setting for a parameter-specific bias diagnostic.
 The random-effects structure involves thousands of teacher effects, the student-level error covariance matrix $\mathbf{R}$ is block-diagonal rather than homoskedastic, and the multiple-membership design encoded by $\mathbf{Z}$} is more complex than a simple random-intercept panel.
-The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
-The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic.
-Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices can be constructed.
-This is implicit in the function call, and the formula argument to GPvam only constructs the fixed-effects design matrix $\mathbf{X}$} and parameter vector $\boldsymbol{\beta}$.
-We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6.
-This is the same data set used by Karl et al.
+The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns internal bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
+The GPvam package (Karl et al., 2012, 2013b), illustrated in Listing 3, fits the CP model and includes a function for the bias diagnostic.
+The supplied R data object must contain columns named “teacher”, “year”, and “student”, because those variables determine the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices.
+This construction is implicit in the function call: the formula argument to GPvam specifies only the fixed-effects design matrix $\mathbf{X}$} and coefficient vector $\boldsymbol{\beta}$.
+We illustrate the method with mathematics test scores from a large urban elementary-school dataset containing 2834 students in grades 4 through 6, the same dataset analyzed by Karl et al.
 (2013a).
-In this value-added model application, we use the GPvam implementation of the bias diagnostic, which accommodates the block-diagonal error covariance matrix and multiple-membership random effects structure that fall outside the scope of the current mixedbiastest implementation.
+In this application we use the GPvam implementation of the bias diagnostic because it accommodates the block-diagonal error covariance matrix and multiple-membership random-effects structure that lie outside the scope of the current mixedbiastest implementation.
 1library(GPvam)
@@
 4 hessian = TRUE,tol1=1e-9)
-5# Perform the bias test on all fixed effects
+5# Perform the bias diagnostic on all fixed effects
 6bias.test.custom(result.cp)
-7# Perform the bias test on a contrast: White - Hispanic
+7# Perform the bias diagnostic on a contrast: White - Hispanic
 8bias.test.custom(result.cp,n_perms=1e6, k_vectors = list(c(0,0,0,-1,1,0,0,0)))
 Listing 3:  Fitting a CP model with GPvam and performing the bias diagnostic
-Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Bias p-values
-Fixed Effect Estimate Std. Error Bias Estimate Bias p-value
+Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Permutation $p$-values
+Fixed Effect Estimate Std. Error Bias Estimate Perm. $p$-value
 Black 24.2894 0.1612 0.0062 0.7430
@@
 Gender=M 0.0552 0.0819 0.0036 0.4990
-The CP estimates and bias diagnostics for the fixed effects appear in Table 2.
-The diagnostic indicates that the estimator for the Hispanic effect is downward biased, while the estimators for the White and Asian/Pac Island effects are upward biased.
-The diagnostic for the contrast White - Hispanic reveals even stronger bias: the estimated contrast is $0.1287$, and none of the one million permutations exceeded the test statistic ($p$-value $\approx 0$).
-Although the bias is statistically significant, its magnitude may not be practically significant for this particular application.
-Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis.
-In this example we do not propose a change in estimator based solely on the diagnostic.
-Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.
+The CP estimates and fixed-effect bias diagnostics appear in Table 2.
+The diagnostic suggests downward bias for the Hispanic coefficient and upward bias for the White and Asian/Pac Island coefficients.
+For the contrast White - Hispanic, the internal bias estimate is $0.1287$, and none of the one million sampled permutations produced a value as extreme, so the Monte Carlo $p$-value is effectively 0 at that simulation resolution.
+Although the statistical evidence for bias is strong, its practical importance depends on the scale on which the contrast will be interpreted.
+Even so, the example shows how the diagnostic can be used to investigate possible bias arising from nonrandom student assignment on a coefficient-by-coefficient basis.
+We do not propose changing estimators on the basis of the diagnostic alone; instead, we use the bias estimates descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.
 5. Conclusion
-We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects.
-Besides considering bias rather than consistency, this diagnostic is narrower than the Hausman test in the sense that it targets the component of bias captured by $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$, i.e., bias operating through systematic alignment of the random effects with the weighting vector $\hat{\boldsymbol{\nu_{k}}}$ when the fixed- and random-effects design matrices are not orthogonal (with respect to $\mathbf{R}^{-1}$).
-In the stochastic-$\mathbf{Z}$} setting emphasized by Karl & Zimmerman (2021), such alignment is induced by dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$.
-In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent.
-Nevertheless, the bias diagnostic can be computed from the fitted random-effects model without changing the original RE specification and generates summaries for each fixed effect or contrast.
-These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.
+We have illustrated how the diagnostic of Karl & Zimmerman (2021) can be used to flag fixed effects and contrasts for which a fitted mixed model may exhibit non-negligible finite-sample bias.
+Relative to the Hausman test, the diagnostic is narrower: it targets the coefficient-specific alignment term $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$ rather than every possible source of inconsistency in the RE estimator.
+Its formal sampling interpretation is sharpest in the stochastic-$\mathbf{Z}$} settings emphasized by Karl & Zimmerman (2021), where nonrandom assignment can induce alignment between the random-effects design and the random effects themselves.
+In more conventional panel settings, we view the resulting statistic primarily as a descriptive complement to FE–RE comparisons.
+Even so, the diagnostic is attractive because it can be computed from a single fitted RE model and reported for any fixed effect or contrast of substantive interest.
 In applications where both RE and FE (or CRE) estimators are readily available, we view the Hausman or Mundlak–Wooldridge tests as primary specification checks: a rejection of the null typically leads the analyst to prefer an estimator that is robust to correlation between regressors and random effects.